OpenStudy (anonymous):

Write the slope intercept equation for the line with slope -1/4 and y-intercept (0,-2).

OpenStudy (cwrw238):

y - y1 = m(x - x1) y - -2 = (-1/4) x y = (-1/4)x - 2

OpenStudy (saifoo.khan):

\[y = -\frac14x -2\]

OpenStudy (anonymous):

thank you guys! I had a brain fart I am so exhausted and could not think:)

OpenStudy (saifoo.khan):

You welcome!